# ECG-Chat: A Large ECG-Language Model for Cardiac Disease Diagnosis

### v1

@@ Appendix F Appendix F: Ablation Study
Table 6: Performance comparison between single dataset and mixed dataset.
@@
Table 7: Performance comparison between different text encoders.
Text Encoder. We selected three different pretrained text encoders, Med-CPT (Jin et al. 2023) obtained on the contrastive learning task, and BioLinkBert (Yasunaga, Leskovec, and Liang 2022) and BioClinicalBert (Alsentzer et al. 2019) obtained on the reconstruction task. As shown in Table 7, BioLinkBert and Med-CPT achieved the best results in retrieval and classification tasks respectively.
Table 8: Scalability analysis results
@@
Scalability. We train the model on 10%, 50% and 100% of the data, as shown in Table 8(a). Using all the data, we train models with 43M, 85M and 128M parameters, respectively, by changing the number of transformer layers. The results are shown in Table 8(b). It can be seen that the number of model parameters has little impact on the results. When increasing the amount of training data, the model effect can be significantly improved, especially on retrieval tasks.
Appendix G Appendix G: Case Study

### v2

@@ Appendix F Ablation Study
TABLE IV: Performance comparison between single dataset and mixed dataset.
@@
TABLE V: Performance comparison between different text encoders.
Text Encoder
We selected three different pretrained text encoders, Med-CPT [34] obtained on the contrastive learning task, and BioLinkBert [52] and BioClinicalBert [53] obtained on the reconstruction task. As shown in Table V, BioLinkBert and Med-CPT achieved the best results in retrieval and classification tasks respectively.
TABLE VI: Scalability analysis results
@@
Scalability
We train the model on 10%, 50% and 100% of the data, as shown in Table VI(a). Using all the data, we train models with 43M, 85M and 128M parameters, respectively, by changing the number of transformer layers. The results are shown in Table VI(b). It can be seen that the number of model parameters has little impact on the results. When increasing the amount of training data, the model effect can be significantly improved, especially on retrieval tasks.
GraphRAG DSPy F AR CR CP CU CER SS
39.87 32.56 9.03 29.82 29.13 24.67 18.94
76.60 68.29 32.67 67.53 64.39 57.00 54.40
71.57 63.93 27.57 62.81 59.54 53.35 47.72
82.12 74.29 39.44 73.18 88.92 73.10 81.83
TABLE VII: Ablation Study of GraphRAG and DSPy in Model Performance Metrics
Effectiveness of GraphRAG and DSPy
Table VII shows the combination of GraphRAG and DSPy significantly outperforms other combinations across all metrics. Without GraphRAG and DSPy, the model’s performance is relatively poor, with a Faithfulness score of 39.87 and Answer Relevancy score of 32.56, among others. Introducing the GraphRAG module results in a significant improvement in Faithfulness (76.60) and Answer Relevancy (68.29). Meanwhile, incorporating the DSPy module leads to improvements in Context Recall (27.57) and Summarization Score (47.72).
Appendix G Case Study